OpenStudy (anonymous):

one more problem i been stuck on find equations of the lines that bisect the acute and obtuse angles formed by the graphs of 6x + 8y = -5 and 2x - 3y = 4

OpenStudy (anonymous):

Take a point (u,v) and fine a relation between u and v so the distance from (u,v) to the two lines are the same

OpenStudy (anonymous):

Square the two distances and make them equal, you get \[ \frac{(6 u+8 v+5)^2}{6^2+8^2}=\frac{(2 u-3 v-4)^2}{2^2+3^2} \]

OpenStudy (anonymous):

Solve for v in terms of u

OpenStudy (anonymous):

so wait how do i find the point between u and v?

OpenStudy (anonymous):

Take any point (u,v). If it is on the bisector the distance from (u,v) to each of the line must be the same

OpenStudy (anonymous):

so it is okay if I can use 2 and 1

OpenStudy (anonymous):

You will get the two bisectors at once \[ v=\left(18-5 \sqrt{13}\right) u+\frac{5 \sqrt{13}}{2}-10\\ v=\left(18+5 \sqrt{13}\right) u-\frac{5 \sqrt{13}}{2}-10 \]

OpenStudy (anonymous):

No, (u,v) should be variables. Now exchanging u by x and v by y, you get the equations of the two bissectors \[ y=\left(18-5 \sqrt{13}\right) x+\frac{5 \sqrt{13}}{2}-10\\ y=\left(18+5 \sqrt{13}\right) x-\frac{5 \sqrt{13}}{2}-10 \]

OpenStudy (anonymous):

Here is the graph of the lines and the bisectors

OpenStudy (anonymous):

ohh wow thank you im taking my time to understand this concept

OpenStudy (anonymous):

Notice that the product of the two slopes of the bisectors is =-1. Did we expect that? \[ \left(18-5 \sqrt{13}\right) \left(18+5 \sqrt{13}\right)=-1 \]

OpenStudy (anonymous):

The concept is easy, working out the solution is a bit tedious, but only requires the use of the quadratic formula.

OpenStudy (anonymous):

ohh thank you very much without your help I would stay up really late just to get this figure out or worst I just skipped the problem

OpenStudy (anonymous):

YW